OpenStudy (anonymous):

Please help!! Dx Solve sin(x)(sin(x)-1)=0

OpenStudy (anonymous):

sin x = 0 or sin x = 1 x = 0,180,360 or x = 90,270

OpenStudy (anonymous):

x = 90 only not include 270, sorry

OpenStudy (anonymous):

\[ \sin(x) =0 => x = 2 k \pi, \,\,k=0, \pm 1, \pm2 \cdots \] \[ \sin(x) =1 => x = \frac \pi 2+2 k \pi, \,\,k=0, \pm 1, \pm2 \cdots \]

OpenStudy (anonymous):

Choose the k above to fit your interval for x